OpenStudy (anonymous):

find gof if f(x) = x + 7 and g(x) = 3x - 2. A. 3x + 5 B. 3x - 5 C. -3x - 19 D. 3x + 19

OpenStudy (zzr0ck3r):

\((g\circ f)(x)=g(f(x)) = g(x+7)= 3(\color{blue}{x+7})-2\)

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

so which one is it? A,B,C, or D?